OpenStudy (anonymous):

Provide the missing reason in the following proof. A. Definition of Subtraction B. Additive Inverse Property C. Additive Identity Property D. Opposite of a Sum Property

OpenStudy (anonymous):

Prove –(–y – x) – x = y Statements Reasons –(–y – x) – x = –[–y +(–x)] – x Definition of subtraction –[–y +(–x)] – x = y + x – x Opposite of a sum property y + x – x = y + x + (–x) ________________________ y + x + (–x) = y + [x + (–x)] Associative Property of Addition y + [x + (–x)] = y + 0 Additive Inverse Property y + 0 = y Additive Identity Property

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

So what do u want ???

OpenStudy (anonymous):

What do u want ? Whats your problem????

OpenStudy (anonymous):

u have to figure out if its a, b, c, or d for the live that looks like ______________

OpenStudy (anonymous):

and wats the question ???

OpenStudy (anonymous):

why dont u read it

OpenStudy (anonymous):

Provide the missing reason in the following proof.

OpenStudy (anonymous):

...